# Uncertainty Modeling in Graph Neural Networks via Stochastic Differential Equations

### v3

@@ Abstract
We address the problem of learning uncertainty-aware representations for graph-structured data. While Graph Neural Ordinary Differential Equations (GNODE) are effective in learning node representations, they fail to quantify uncertainty. To address this, we introduce Latent Graph Neural Stochastic Differential Equations (LGNSDE), which enhance GNODE by embedding randomness through Brownian motion to quantify uncertainty. We provide theoretical guarantees for LGNSDE and empirically show better performance in uncertainty quantification.
@@ 1 Introduction
Before the widespread of neural networks and the boom in modern machine learning, complex systems in various scientific fields were predominantly modelled using differential equations. Stochastic Differential Equations (SDEs) were the standard approach to incorporating randomness. These methods were foundational across disciplines such as physics, finance, and computational biology (Hoops et al., 2016; Quach et al., 2007; Mandelzweig and Tabakin, 2001; Cardelli, 2008; Buckdahn et al., 2011; Cvijovic et al., 2014).
@@
To address this gap, we propose Latent Graph Neural Stochastic Differential Equations (LGNSDE), a method that perturbs features during both the training and testing phases using Brownian motion noise, allowing for handling noise and aleatoric uncertainty. We also assume a prior SDE latent space and learn a posterior SDE using a GNN. This Bayesian approach to the latent space allows us to quantify epistemic uncertainty. As a result, our model can capture and quantify both epistemic and aleatoric uncertainties. More specifically, our contributions are as follows:
@@
  • We empirically show that Latent GNSDEs demonstrate exceptional performance in uncertainty quantification, outperforming Bayesian GCNs (Hasanzadeh et al., 2020), and GCN ensembles (Lin et al., 2022).
@@ 2 Methodology
Figure 1: The diagram shows the evolution of input data in latent space $\mathbf{H}(t)$ through an SDE, with sample paths (purple) and confidence bands representing variance. At three timesteps, we visualize graph embeddings, where nodes (white and orange) become more separable over time due to the influence of the vector field. The inset axes represent latent dimensions, while the purple and yellow background highlights the magnitude and direction of the vector field guiding the latent dynamics.
@@ 2.1 Model Definition
where $\mathbf{F}_{\mathcal{G}}$ is parameterized by a GCN with $\phi$ representing the learned weights of the neural network. The drift function mainly determines the dynamics of the evolution of the latent state, while the diffusion term $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t))\,\text{d}\mathbf{W}(t)$ introduces stochastic elements. With the need to keep the Kullback-Leibler (KL) divergence bounded, we set the diffusion functions of the prior and posterior to be the same [Calvo-Ordonez et al. 2024; Archambeau et al. 2007].
@@
$\mathcal{L}_{\text{ELBO}}(\phi)=\mathbb{E}\left[\log{p(\mathbf{Y}|\mathbf{H}(t_{1}))}-\int_{t_{0}}^{t_{1}}\frac{1}{2}\left\|v(\mathbf{H}(u),\phi,\theta,\mathcal{G})\right\|_{2}^{2}\,\text{d}u\right],$
@@
To sample $\mathbf{H}(t)$ from the approximate posterior, we integrate the SDE in Eq. 1:
$\mathbf{H}(t_{1})$ $=\mathbf{H}(t_{0})+\int_{t_{0}}^{t_{1}}\mathbf{F}_{\mathcal{G},\phi}(\mathbf{H}(u),u)\,\text{d}u+\int_{t_{0}}^{t_{1}}\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)\,\text{d}\mathbf{W}(u),$
where $\mathbf{H}(t_{0})$ are the node-wise features $\mathbf{X}_{\text{in}}$ in the graph $\mathcal{G}$. In practice, this is not feasible since the posterior drift $\mathbf{F}_{\mathcal{G},\phi}$ is parametrised by a neural network. We numerically solve this integral with a standard Stochastic Runge-Kutta method (Rößler, 2010). We then use a Monte Carlo approximation to get the expectation of $\mathbf{H}(t_{0})$ and approximate the posterior predictive distribution as
$p(\mathbf{Y}^{*}|\mathcal{G},\mathbf{X}_{\text{in}},\mathbf{Y})$ $\approx\frac{1}{N}\sum_{n=1}^{N}p\left(\mathbf{Y}^{*}|\mathbf{H}_{n}(t_{1}),\mathcal{G}\right),$
where $\mathbf{H}_{1}(t_{1}),\ldots,\mathbf{H}_{N}(t_{1})$ are samples drawn from the approximate posterior $p(\mathbf{H}(t_{1})|\mathbf{Y},\mathbf{X}_{\text{in}},\mathcal{G})$.
Following Poli et al. (2019), we use a similar encoder-decoder setup. Our encoding focuses solely on the features of individual nodes, while the graph structure remains unchanged. Finally, we remark that the memory and time complexity are $\mathcal{O}(|\mathcal{E}|d+L)$ and $\mathcal{O}(L)$ respectively, where $L$ is the number of SDE solver steps, $\mathcal{E}$ is the number of edges in the graph and $d$ is the dimension of the features.
@@ 3 Theoretical Guarantees
In this section, we present key results on the stability and robustness of our framework under mild assumptions (Appendix A.2). Firstly, we address the fundamental question of whether our proposed models provide meaningful uncertainties. By showing that the variance of the latent representation bounds the model output variance, we highlight the ability of LGNSDEs to capture and quantify inherent uncertainty in the system. The latent representation is the underlying structure from which the model’s output is generated, i.e. the uncertainty in the latent space directly influences the uncertainty in predictions. We formalize this in the following lemma:
Lemma 1.
Under assumptions 1-3, there exists a unique strong solution to an LGNSDE of the form
$d\mathbf{H}(t)=\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\boldsymbol{\theta})\,dt+\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t),$
whose variance bounds the variance of the model output $\mathbf{y}(t)$ as:
$\text{Var}(\mathbf{y}(t))\leq L^{2}_{h}\text{Var}(\mathbf{H}(t)),$
hence providing a meaningful measure of the total uncertainty.
@@
We now demonstrate the robustness of our framework under small perturbations in the initial conditions. By deriving explicit bounds on the deviation between the perturbed and unperturbed solutions over time, we show that the model’s output remains stable.
Lemma 2.
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time $t$, specifically
@@ 4 Experiments
We evaluate LGNSDEs on 5 datasets (see A.2 for details on these datasets and hyperparameters), we compare it to GNODE (Poli et al., 2019), GCN Kipf and Welling (2016), Bayesian GCN (BGCN) (Hasanzadeh et al., 2020), and an ensemble of GCNs (Lin et al., 2022).
@@
The top figure 2 shows the entropy distributions of the models for correct and incorrect predictions. Note that most models display similar mean entropy for both correct and incorrect predictions. Notably, our model stands out with the largest difference in entropy, with incorrect predictions having 35% more entropy compared to correct predictions, a larger gap than observed in other models.
@@ 4.1 Out of Distribution Detection
LGNSDE (Our) 0.0168 ± 0.0070 0.0479 ± 0.0109 0.0061 ± 0.0011 0.0068 ± 0.0008 0.3205 ± 0.0135
@@
We evaluate the models’ ability to detect out-of-distribution (OOD) data by training them with one class left out of the dataset. This introduces an additional class in the validation and test sets that the models have not encountered during training. The goal is to determine if the models can identify this class as OOD. We analyze the entropy, $H(y|\mathbf{X}_{i})=-\sum_{c=1}^{C}p(y=c|\mathbf{X}_{i})\log p(y=c|\mathbf{X}_{i}),$ where $p(y=c|\mathbf{X}_{i})$ represents the probability of input $\mathbf{X}_{i}$ belonging to class $c$. Entropy quantifies the uncertainty in the model’s predicted probability distribution over $C$ classes for a given input $\mathbf{X}_{i}$.
@@
Figure 2 shows the test entropy distribution for in-distribution (blue) and out-of-distribution (red) data. For each test sample, predictions were made over $C-1$ classes, excluding the left-out class. The OOD class exhibits higher entropy, indicating greater uncertainty. While most models show similar entropy distributions for both data types, our LGNSDE model achieves a clear separation, with a 50% higher mean entropy for OOD data compared to in-distribution data. Other models show less than a 10% difference between the two distributions.

### v4

@@ Abstract
We propose a novel Stochastic Differential Equation (SDE) framework to address the problem of learning uncertainty-aware representations for graph-structured data. While Graph Neural Ordinary Differential Equations (GNODEs) have shown promise in learning node representations, they lack the ability to quantify uncertainty. To address this, we introduce Latent Graph Neural Stochastic Differential Equations (LGNSDE), which enhance GNODE by embedding randomness through a Bayesian prior-posterior mechanism for epistemic uncertainty and Brownian motion for aleatoric uncertainty. By leveraging the existence and uniqueness of solutions to graph-based SDEs, we prove that the variance of the latent space bounds the variance of model outputs, thereby providing theoretically sensible guarantees for the uncertainty estimates. Furthermore, we show mathematically that LGNSDEs are robust to small perturbations in the input, maintaining stability over time. Empirical results across several benchmarks demonstrate that our framework is competitive in out-of-distribution detection, robustness to noise, and active learning, underscoring the ability of LGNSDEs to quantify uncertainty reliably.
@@ 1 Introduction
Before the widespread of neural networks and the boom in modern machine learning, complex systems in various scientific fields were predominantly modelled using differential equations. Stochastic Differential Equations (SDEs) were the standard approach to incorporating randomness. These methods were foundational across disciplines such as physics, finance, and computational biology (Hoops et al., 2016; Quach et al., 2007; Mandelzweig & Tabakin, 2001; Arroyo et al., 2024; Moreno-Pino et al., 2024; Cardelli, 2008; Buckdahn et al., 2011; Cvijovic et al., 2014).
@@
To address this gap, we propose Latent Graph Neural Stochastic Differential Equations (LGNSDE), a method that perturbs features during both the training and testing phases using Brownian motion noise, allowing for handling noise and aleatoric uncertainty. We assume a prior SDE on the latent space and learn a posterior SDE using a GNN as the drift function. This Bayesian approach to the latent space allows us to quantify epistemic uncertainty. As a result, our model can capture and quantify both epistemic and aleatoric uncertainties. More specifically, our contributions are as follows:
@@
  • We empirically show that Latent GNSDEs demonstrate exceptional performance in uncertainty quantification, outperforming Bayesian GCNs (Hasanzadeh et al., 2020), GCN ensembles (Lin et al., 2022) and Graph Gaussian Process (Borovitskiy et al., 2021).
Figure 1: The diagram shows the evolution of one of the nodes of the input graph in latent space, $\mathbf{H}(t)$, through an SDE, with sample paths (purple) and confidence bands representing variance. At three timesteps, we visualize graph embeddings, where nodes (white and orange) become more separable over time due to the influence of the vector field. The inset axes represent latent dimensions, while the purple and yellow background highlights the magnitude and direction of the vector field guiding the latent dynamics.
2 Background
Graph Neural Ordinary Differential Equations (GNODE).
Introduced by Poli et al. (2019), GNODEs extend Graph ResNets by modelling node representations continuously over time. In a typical Graph ResNet, the node features evolve according to the update rule
$\mathbf{H}(t+1)=\mathbf{H}(t)+\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\theta),$
where $t$ represents the layer index, $\mathbf{H}(t)$ and $\mathbf{H}(t+1)$ are the input and output node embeddings, and $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$. The function $f_{\theta}$, parameterized by $\theta$, defines the transformation applied to node features. Here, the input features are $\mathbf{X}_{\text{in}}=\mathbf{H}(0)$, and the final node embeddings $\mathbf{H}(T)$ yield the model’s output predictions, denoted $\mathbf{\hat{Y}}$. Now, consider an update with a small time step $c\in\mathbb{R}$
$\mathbf{H}(t+c)=\mathbf{H}(t)+c\cdot\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\theta),$
which leads to the continuous limit as $c\to 0$
$\frac{\text{d}\mathbf{H}(t)}{\text{d}t}=\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\theta).$
This differential equation represents the continuous evolution of node features over time, transforming the discrete depth of layers into a continuous variable $t$. The solution to this equation is given by
$\mathbf{H}(t)=\mathbf{H}(0)+\int_{0}^{t}\mathbf{F}_{\mathcal{G}}(\mathbf{H}(u),u,\theta)\,\text{d}u.$
In practice, $\mathbf{F}_{\mathcal{G}}$ is modelled by a neural network, and $t$ operates as a continuous depth parameter. The solution to the integral is approximated numerically, making GNODE a continuous-depth analogue of graph residual networks.
@@ 3.1 Model Definition
where $\mathbf{F}_{\mathcal{G}}$ is now parameterized by a GCN with $\phi$ representing the learned weights of the neural network. The drift function mainly determines the dynamics of the evolution of the latent state, while the diffusion term $\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t))\,\text{d}\mathbf{W}(t)$ introduces stochastic elements. With the need to keep the Kullback-Leibler (KL) divergence bounded, we set the diffusion functions, $\mathbf{G}_{\mathcal{G}}$, of the prior and posterior to be the same [Calvo-Ordonez et al. 2024; Archambeau et al. 2007].
@@
$\mathcal{L}_{\text{ELBO}}(\phi)=\mathbb{E}\left[\log{p(\mathbf{Y}|\mathbf{H}(t))}-\int_{0}^{t}\frac{1}{2}\left\|v(\mathbf{H}(u),\phi,\theta,\mathcal{G})\right\|_{2}^{2}\,\text{d}u\right],$
@@
We sample $\mathbf{H}(t)$ by integrating the SDE in Eq. 1. The analytical solution is
$\mathbf{H}(t)$ $=\mathbf{H}(0)+\int_{0}^{t}\mathbf{F}_{\mathcal{G},\phi}(\mathbf{H}(u),u)\,\text{d}u+\int_{0}^{t}\mathbf{G}_{\mathcal{G}}(\mathbf{H}(u),u)\,\text{d}\mathbf{W}(u),$
where $\mathbf{H}(0)$ are the node-wise features $\mathbf{X}_{\text{in}}$ in the graph $\mathcal{G}$. We numerically solve this integral with a standard Stochastic Runge-Kutta method (Rößler, 2010). We then use a Monte Carlo approximation to get the expectation of $\mathbf{H}(t)$ and approximate the posterior predictive distribution as
$p(\mathbf{Y}^{*}|\mathcal{G},\mathbf{X}_{\text{in}},\mathbf{Y})$ $\approx\frac{1}{N}\sum_{n=1}^{N}p\left(\mathbf{Y}^{*}|\mathbf{H}_{n}(t),\mathcal{G}\right),$
where $\mathbf{H}_{1}(t_{1}),\ldots,\mathbf{H}_{N}(t)$ are samples drawn from the approximation to $p(\mathbf{H}(t)|\mathbf{Y},\mathbf{X}_{\text{in}},\mathcal{G})$.
Following Poli et al. (2019), we use a similar encoder-decoder setup. Our encoding focuses solely on the features of individual nodes, while the graph structure remains unchanged. Finally, we remark that the memory complexity when using the stochastic adjoint sensitivity method is $\mathcal{O}(1)$ and the time complexity is $\mathcal{O}(L\log L(|\mathcal{E}|d+|\mathcal{V}|d))$, where $L$ is the number of SDE solver steps, $\mathcal{E}$ is the number of edges in the graph, $\mathcal{V}$ is the number of nodes, and $d$ is the dimension of the features (see Appendix C.3). For a runtime comparison with other models see 12, and 15.
Note that in our framework, model depth is inherently tied to the evolution of the latent space, where depth is determined by the number of layers corresponding to the sampling steps of the SDE solver. As the SDE solver dictates the number of steps, it effectively controls the number of layers in the model. Thereby the SDE solver chooses the number of layers, dynamically changing the model’s complexity based on task difficulty. For more complex tasks, the solver will generate additional steps (layers), while simpler tasks will require fewer layers. Exploring optimal adaptive SDE solvers will remain part of future work.
@@ 4 Theoretical Guarantees
To establish the theoretical foundations of our framework, we begin by acknowledging the existence and uniqueness results for graph-based stochastic differential equations, as proven in Lin et al. (2024). This result guarantees that, under certain assumptions (Appendix A.2), there exists a unique mild solution to the Graph Neural SDE, ensuring the well-posedness of the model’s dynamics and that the solution behaves in a stable and predictable manner, i.e. small changes in the initial conditions or input data lead to small changes in the solution. We borrow this result in the following theorem:
Theorem 1 (Lin et al. (2024)).
If $\Psi$ is a linear operator with a complete orthonormal basis set and eigenvalues $\lambda_{k}>0$, the continuous operator $\mathbf{G}_{\mathcal{G}}$ satisfies the Lipschitz condition, and the initial node representation $\mathbf{H}_{i}(0)$ is square-integrable and $\mathcal{F}_{0}$-measurable, then there exists a unique mild solution $\mathbf{H}_{i}(t)$ on $[0,T]$ for any $T>0$ and $i\in\mathcal{V}$, such that
[…]
where $e^{t\Psi}$ is the semigroup generated by $\Psi$. Furthermore, there exists a constant $C_{T}>0$ such that
$\sup_{t\in[0,T]}\|\mathbf{H}_{i}(t)\|\leq C_{T}(1+\|\mathbf{H}_{i}(0)\|).$
This theorem confirms that graph-based SDEs have a well-posed solution trajectory over time that is well-behaved and bounded, ensuring that our LGNSDE model can maintain stability across varying graph structures, meaning that the solution does not exhibit erratic or unbounded behaviour even under small changes in graph structure or input features.
Leveraging this result, we proceed to present key results on the stability and robustness of our framework.
  • We derive a bound that proves that our proposed models provide meaningful uncertainties.
  • We demonstrate the robustness of our framework under small perturbations in the initial conditions.
@@
By showing that the variance of the latent representation bounds the model output variance, we highlight the ability of LGNSDEs to capture and quantify inherent uncertainty in the system. The model’s output is given by the trajectory of the latent representation since $\mathbf{y}=\mathbf{H}(t=T)$, therefore the uncertainty in the latent space directly influences the uncertainty in predictions. We formalize this in the following proposition:
Proposition 1.
Under assumptions 1-3 and given Theorem 1, there exists a unique mild [^1] A mild solution to an SDE is expressed via an integral equation involving the semigroup generated by the linear operator and represents a weaker notion of the solution. solution to an LGNSDE of the form
$d\mathbf{H}(t)=\mathbf{F}_{\mathcal{G}}(\mathbf{H}(t),t,\bm{\theta})\,dt+\mathbf{G}_{\mathcal{G}}(\mathbf{H}(t),t)\,d\mathbf{W}(t),$
whose variance bounds the variance of the model output $\mathbf{\hat{y}}(t)$ as:
$\text{Var}(\mathbf{\hat{y}}(t))\leq L^{2}_{h}\text{Var}(\mathbf{H}(t)),$
where $L^{2}_{h}$ is the Lipschitz constant of the readout layer. This ensures that the output variance is bounded by the prior variance of the latent space, providing a controlled measure of uncertainty.
@@
Now, by deriving explicit bounds on the deviation between the perturbed and unperturbed solutions over time, we show that the model’s output remains stable.
Proposition 2.
@@
Under assumptions 1-3, consider two initial conditions $\mathbf{H}_{0}$ and $\tilde{\mathbf{H}}_{0}=\mathbf{H}_{0}+\delta\mathbf{H}(0)$, where $\delta\mathbf{H}(0)\in\mathbb{R}^{n\times d}$ is a small perturbation in the initial node features with $\|\delta\mathbf{H}(0)\|_{F}=\epsilon$. Assume that $\mathbf{H}_{0}$ is taken from a compact set $\mathcal{H}\subseteq\mathbb{R}^{n\times d}$. Then, the deviation between the solutions $\mathbf{H}(t)$ and $\tilde{\mathbf{H}}(t)$ of the LGNSDE with these initial conditions remains bounded across time $t$ [^2] Note that while the bound is exponential in $t$, in practice, the time horizon is usually constrained to a limited range, such as $t\in[0,1]$. Within this interval, the exponential factor does not grow excessively, ensuring that the deviation between the perturbed and unperturbed solutions remains under control., specifically
@@ 5 Experiments
Metric Model Cora Citeseer Computers Photo Pubmed
AUROC (↑) GNN 0.9654 ± 0.0050 0.9173 ± 0.0068 0.9680 ± 0.0016 0.9905 ± 0.0003 0.9006 ± 0.0139
GNODE 0.9664 ± 0.0051 0.9146 ± 0.0063 0.9569 ± 0.0067 0.9885 ± 0.0007 0.8857 ± 0.0203
GRAPH GP 0.8970 ± 0.0055 0.8877 ± 0.0062 OOM OOM OOM
LGNSDE (Ours) 0.9659 ± 0.0038 0.9111 ± 0.0072 0.9691 ± 0.0032 0.9909 ± 0.0004 0.9004 ± 0.0087
AURC (↓) GNN 0.0709 ± 0.0101 0.1626 ± 0.0109 0.0745 ± 0.0053 0.0199 ± 0.0007 0.1367 ± 0.0192
GNODE 0.0628 ± 0.0095 0.1609 ± 0.0141 0.1055 ± 0.0165 0.0219 ± 0.0015 0.1427 ± 0.0214
BGCN 0.0858 ± 0.0219 0.1764 ± 0.0215 0.1634 ± 0.0344 0.1218 ± 0.0577 0.4152 ± 0.1723
ENSEMBLE 0.0789 ± 0.0061 0.1722 ± 0.0179 0.0877 ± 0.0037 0.0244 ± 0.0012 0.1722 ± 0.0285
GRAPH GP 0.1869 ± 0.0084 0.2328 ± 0.0118 OOM OOM OOM
LGNSDE (Ours) 0.0702 ± 0.0095 0.1686 ± 0.0146 0.0687 ± 0.0114 0.0186 ± 0.0007 0.1378 ± 0.0118
Accuracy (↑) GNN 0.8105 ± 0.0173 0.7258 ± 0.0137 0.8098 ± 0.0048 0.9116 ± 0.0021 0.7570 ± 0.0229
GNODE 0.8202 ± 0.0149 0.7235 ± 0.0159 0.7911 ± 0.0098 0.9053 ± 0.0032 0.7577 ± 0.0231
GRAPH GP 0.6491 ± 0.0116 0.6675 ± 0.0127 OOM OOM OOM
LGNSDE (Ours) 0.8079 ± 0.0154 0.7120 ± 0.0119 0.8247 ± 0.0103 0.9169 ± 0.0021 0.7589 ± 0.0161
Table 1: Performance comparison of models across five datasets (Cora, Citeseer, Computers, Photo, Pubmed) based on AUROC, AURC, and Accuracy (mean ± standard deviation). Red indicates the best-performing model, while blue indicates the second-best-performing model for each metric. Some results for the Graph GP model are unavailable due to out-of-memory (OOM) errors.
We evaluate LGNSDE on the following datasets: Cora (Sen et al., 2008), CiteSeer (Giles et al., 1998), PubMed (Sen et al., 2008), and the Amazon co-purchasing graphs Computer (McAuley et al., 2015) and Photo (Shchur et al., 2018). We compare its performance against GNODE (Poli et al., 2019), GCN (Kipf & Welling, 2016), Bayesian GCN (BGCN) (Hasanzadeh et al., 2020), an ensemble of GCNs (Lin et al., 2022) and Graph GPs [^3] We were unable to run the Graph GP model on certain datasets due to its high memory requirements and scalability issues, resulting in out-of-memory (OOM) errors. Consequently, we report these cases as OOM in our experimental results tables. (Borovitskiy et al., 2021).
In conducting our experiments, we used the setup outlined in Shchur et al. (2018). This involved using 20 random weight initializations for datasets with fixed Planetoid splits and implementing 100 random splits for other datasets. The hyperparameters that achieved the highest validation accuracy were chosen, and their performance was evaluated on a test set. For further details on our hyperparameter grid search refer to Appendix B.
5.1 Standard Setting
@@
The top of Figure 2 shows the entropy distributions of the models for correct and incorrect predictions. Note that most models display similar mean entropy for both correct and incorrect predictions. Notably, our model stands out with the largest difference in entropy, with incorrect predictions having 35% more entropy (more uncertainty – see Subsection 5.2) compared to correct predictions, a larger gap than observed in other models.
@@ 5.2 Out of Distribution Detection
LGNSDE (Ours) 0.0168 ± 0.0070 0.0479 ± 0.0109 0.0061 ± 0.0011 0.0068 ± 0.0008 0.3205 ± 0.0135
Accuracy (↑) GNN 0.9470 ± 0.0004 0.8614 ± 0.0071 0.9788 ± 0.0000 0.9558 ± 0.0002 0.6180 ± 0.0155
GNODE 0.9469 ± 0.0002 0.8603 ± 0.0086 0.9788 ± 0.0004 0.9557 ± 0.0000 0.6084 ± 0.0120
BGCN 0.9472 ± 0.0004 0.8711 ± 0.0133 0.9797 ± 0.0010 0.9558 ± 0.0002 0.6039 ± 0.0074
ENSEMBLE 0.9470 ± 0.0003 0.8699 ± 0.0113 0.9788 ± 0.0001 0.9560 ± 0.0001 0.6216 ± 0.0112
LGNSDE (Ours) 0.9471 ± 0.0003 0.8729 ± 0.0108 0.9788 ± 0.0000 0.9560 ± 0.0002 0.6243 ± 0.0094
Table 2: Performance comparison of models for OOD detection across five datasets (Cora, Citeseer, Computers, Photo, Pubmed). Metrics reported are AUROC, AURC, and Accuracy (mean ± standard deviation). Red indicates the best-performing model, while blue indicates the second-best-performing model for each metric.
@@
We evaluate the models’ ability to detect out-of-distribution (OOD) data by training them with one class left out of the dataset. This introduces an additional class in the validation and test sets that the models have not encountered during training. The goal is to determine if the models can identify this class as OOD. We analyze the entropy
$H(\hat{y}|\mathbf{X}_{i})=-\sum_{c=1}^{C}p(\hat{y}=c|\mathbf{X}_{i})\log p(\hat{y}|\mathbf{X}_{i}),$ (2)
where $p(\hat{y}=c|\mathbf{X}_{i})$ represents the probability of input $\mathbf{X}_{i}$ belonging to class $c$. Entropy quantifies the uncertainty in the model’s predicted probability distribution over $C$ classes for a given input $\mathbf{X}_{i}$.
@@
The bottom of Figure 2 shows the test entropy distribution for in-distribution (blue) and out-of-distribution (red) data. For each test sample, predictions were made over $C-1$ classes, excluding the left-out class. The OOD class exhibits higher entropy, indicating greater uncertainty. While most models show similar entropy distributions for both data types, our LGNSDE model achieves a clear separation, with a 50% higher mean entropy for OOD data compared to in-distribution data. Other models show less than a 10% difference between the two distributions.